### v4

@@ 1.5.2 Online Calibration Algorithms
Recent research has made significant progress in proving upper and lower bounds on the optimal rate achievable for both ECE and the distance to calibration in online binary prediction, though significant gaps remain between the current best upper and lower bounds. For ECE minimization, Foster and Vohra (1998) design a randomized algorithm that achieves $O(T^{-2/3})$ expected ECE, which remains the best known upper bound. Existence proofs and constructions of such algorithms have been further explored in several subsequent works (Hart, 2022; Foster and Hart, 2021). A recent work (Qiao and Valiant, 2021) show a lower bound of $\Omega(T^{-0.472})$ to online ECE minimization. For the distance to calibration, Qiao and Zheng (2024) give a non-constructive minimax-based proof for an $O(\frac{1}{\sqrt{T}})$ upper bound and an $\Omega(T^{-1/3})$ lower bound for the same problem. Soon afterwards, Arunachaleswaran et al. (2024) provide a construction of an efficient algorithm that achieves the $O(\frac{1}{\sqrt{T}})$ upper bound for the distance to calibration.
@@ 2.4 Calibration Fixed Decision Loss
$\textsc{PSwap}_{S}(\bm{p},\bm{\theta})=\max_{\sigma:\Delta(\Theta)\to A}\frac{1}{T}\sum_{t\in[T]}\bigg{[}S(\sigma(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]},$
where $a^{*}$ is the best response to prediction as defined in Equation 10.
@@
ECE cannot be represented by the CFDL with any proper scoring rule
@@ 3 Calibration Decision Loss
$\textsc{CDL}(\bm{p},\bm{\theta})=\sup_{S\in[0,1]}\textsc{CFDL}_{S}(\bm{p},\bm{\theta}),$
@@ 3.1 Bregman Divergence Formulation
Specified by a convex function $u:[0,1]\to\mathbb{R}$ and its sub-gradient $\nabla u:[0,1]\to\mathbb{R}$, the Bregman divergence $\textsc{Breg}(p^{\prime},p)$ is defined as [^9] The definition of Bregman divergence here reverses the order of input in the conventional definition to align with proper scoring rules. In the conventional definition, $\textsc{Breg}(p,\widehat{p})=u(p)-u(\widehat{p})+\nabla u(\widehat{p})(\widehat{p}-p)$.
@@ 3.2 Approximation via V-Bregman Divergences
$\textsc{Breg}_{S}(p,\widehat{p})=\int_{\mu=0}^{1}u^{\prime\prime}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}\text{d}\mu.$ (16)
Definition 3.9 (VCFDL).
Given a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ from a finite set $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ and realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states, the VCFDL is defined as the supremem over all V-Bregman divergences:
$\textsc{VCFDL}=\sup_{\mu\in[0,1]}\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i}).$ (17)
@@
VCFDL is a constant-factor approximation of CDL. That is, for any sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ and any sequence of states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$,
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
@@
Since $\textsc{VBreg}_{\mu}$ is a special case of a Bregman divergence $\textsc{Breg}_{S}$ for a scoring rule $S$, comparing the definition of VCFDL in (17) and the expression of PSwap in (13), we immediately get $\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})$.
@@
Let us consider any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$. By Theorem 3.4, there exists a convex function $u:[0,1]\to[0,1]$ with bounded sub-gradients $\nabla u(p)\in[-1,1]$ such that (11) holds. We can construct a sequence of twice-differentiable convex functions $u_{1},u_{2},\ldots$ to approximate $u$ in the following manner. For every $q\in\{0,1,q_{1},\ldots,q_{m},\widehat{q}_{1},\ldots,\widehat{q}_{m}\}$, the derivative $u_{j}^{\prime}(q)$ satisfies $u_{j}^{\prime}(q)=\nabla u(q)$ for every $j$, and the function value $u_{j}$ itself satisfies $u_{j}(q)\to u(q)$ as $j\to+\infty$. Let $S_{j}$ denote the proper scoring rule generated by $u_{j}$ as in (11). For every $i=1,\ldots,m$, we have
@@
$\sum_{i\in[m]}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$ $=\sum_{i\in[m]}n_{i}\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$=\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$\leq\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VCFDL}(\bm{p},\bm{\theta})\text{d}\mu$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\text{d}\mu$
$\leq\textsc{VCFDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\text{d}\mu$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))$
$=\textsc{VCFDL}(\bm{p},\bm{\theta})(\nabla u(1)-\nabla u(0))$
$\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
@@
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i})$ $=\lim_{j\to\infty}\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$
$\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta}).$
Taking supremum over $S$ proves $\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCFDL}(\bm{p},\bm{\theta})$. ∎
@@ 4 CDL and Calibration Errors
    By Example 4.2, $\textsc{CDL}\leq 2\textsc{VCFDL}\leq 4\epsilon$. We can find a Bregman divergence such that $\textsc{CFDL}\geq\frac{\epsilon}{2}$. Consider the scoring rule $S$ that has a V-Bregman divergence with kink $\mu=1-\frac{\epsilon}{2}$.
@@
    We can prove $\textsc{VCFDL}=\Theta(1)$. Fix any V-Bregman divergence with kink $\mu$, consider the corresponding $\textsc{CFDL}_{\mu}$.
@@
    We know $\textsc{CDL}\leq 2\textsc{VCFDL}\leq\frac{8}{T}$.
@@ 5.1 Attributing CDL to Bucket-wise Biases
$\textsc{VCFDL}(\bm{p},\bm{\theta})\leq\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$ (19)

### v5

@@ 1.5.2 Online Calibration Algorithms
Recent research has made significant progress in proving upper and lower bounds on the optimal rate achievable for both ECE and the distance to calibration in online binary prediction, though significant gaps remain between the current best upper and lower bounds. For ECE minimization, Foster and Vohra (1998) shows there exists a randomized algorithm that achieves $O(T^{-1/3})$ expected ECE, which remains the best known upper bound. Existence proofs and constructions of such algorithms have been further explored in several subsequent works (Hart, 2022; Foster and Hart, 2021). A recent work (Qiao and Valiant, 2021) show a lower bound of $\Omega(T^{-0.472})$ to online ECE minimization. For the distance to calibration, Qiao and Zheng (2024) give a non-constructive minimax-based proof for an $O(\frac{1}{\sqrt{T}})$ upper bound and an $\Omega(T^{-1/3})$ lower bound for the same problem. Soon afterwards, Arunachaleswaran et al. (2024) provide a construction of an efficient algorithm that achieves the $O(\frac{1}{\sqrt{T}})$ upper bound for the distance to calibration.
@@ 2.4 Calibration Fixed Decision Loss
$\textsc{PSwap}_{S}(\bm{p},\bm{\theta})=\max_{\sigma:\Delta(\Theta)\to\Delta(\Theta)}\frac{1}{T}\sum_{t\in[T]}\bigg{[}S(\sigma(p_{t}),\theta_{t})-S(p_{t},\theta_{t})\bigg{]}.$
@@
ECE cannot be represented by the CFDL with any proper scoring rule:
@@ 3 Calibration Decision Loss
$\textsc{CDL}(\bm{p},\bm{\theta})=\sup_{S\in[0,1]}\textsc{CFDL}_{S}(\bm{p},\bm{\theta}),$ (11)
@@ 3.1 Bregman Divergence Formulation
Specified by a convex function $u:[0,1]\to\mathbb{R}$ and its sub-gradient $\nabla u:[0,1]\to\mathbb{R}$, the Bregman divergence $\textsc{Breg}(p,\widehat{p})$ is defined as [^9] The definition of Bregman divergence here reverses the order of input in the conventional definition to align with proper scoring rules. In the conventional definition, $\textsc{Breg}(p,\widehat{p})=u(p)-u(\widehat{p})+\nabla u(\widehat{p})(\widehat{p}-p)$.
@@ 3.2 Approximation via V-Bregman Divergences
$\textsc{Breg}_{S}(p,\widehat{p})=\int_{\mu=0}^{1}u^{\prime\prime}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}(p,\widehat{p})\text{d}\mu.$ (17)
Definition 3.9 (VCDL).
Consider a sequence of $T$ predictions $\bm{p}=(p_{t})_{t\in[T]}$ from a finite set $Q=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$ and realizations $\bm{\theta}=(\theta_{t})_{t\in[T]}$ of states. For $\mu\in[0,1]$, we define $\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ to be the Calibration Fixed Decision Loss for the V-shaped scoring rule $S_{\mu}$ with kink $\mu$, or equivalently, by Proposition 3.6, $\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ can be expressed using the V-Bregman divergence at kink $\mu$:
$\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i}).$
We define $\textsc{VCDL}(\bm{p},\bm{\theta})$ to be the supremum of $\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta})$ over all V-Bregman divergences:
$\textsc{VCDL}(\bm{p},\bm{\theta})=\sup_{\mu\in[0,1]}\textsc{CFDL}_{\mu}(\bm{p},\bm{\theta}).$ (18)
@@
VCDL is a constant-factor approximation of CDL. That is, for any sequence of predictions $\bm{p}=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ and any sequence of states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$,
$\textsc{VCDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCDL}(\bm{p},\bm{\theta}).$
@@
Since $\textsc{VBreg}_{\mu}$ is a special case of a Bregman divergence $\textsc{Breg}_{S}$ for a scoring rule $S$, comparing the definition of VCDL in (18) and the definition of CDL in (11), we immediately get $\textsc{VCDL}(\bm{p},\bm{\theta})\leq\textsc{CDL}(\bm{p},\bm{\theta})$.
@@
Let us consider any proper scoring rule $S:[0,1]\times\{0,1\}\to[0,1]$. By Theorem 3.4, there exists a convex function $u:[0,1]\to[0,1]$ with bounded sub-gradients $\nabla u(p)\in[-1,1]$ such that (12) holds. We can construct a sequence of twice-differentiable convex functions $u_{1},u_{2},\ldots$ on the domain $[0,1]$ to approximate $u$ in the following manner. For every $q\in\{0,1,q_{1},\ldots,q_{m},\widehat{q}_{1},\ldots,\widehat{q}_{m}\}$, the derivative $u_{j}^{\prime}(q)$ satisfies $u_{j}^{\prime}(q)=\nabla u(q)$ for every $j$, and the function value $u_{j}$ itself satisfies $u_{j}(q)\to u(q)$ as $j\to+\infty$. Let $S_{j}$ denote the proper scoring rule generated by $u_{j}$ as in (12). For every $i=1,\ldots,m$, we have
@@
$\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$ $=\frac{1}{T}\sum_{i\in[m]}n_{i}\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$=\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\frac{1}{T}\sum_{i\in[m]}n_{i}\textsc{VBreg}_{\mu}(q_{i},\widehat{q}_{i})\text{d}\mu$
$\leq\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\textsc{VCDL}(\bm{p},\bm{\theta})\text{d}\mu$
$=\textsc{VCDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\max(1-\mu,\mu)\text{d}\mu$
$\leq\textsc{VCDL}(\bm{p},\bm{\theta})\int_{\mu=0}^{1}u^{\prime\prime}_{j}(\mu)\cdot\text{d}\mu$
$=\textsc{VCDL}(\bm{p},\bm{\theta})(u_{j}^{\prime}(1)-u_{j}^{\prime}(0))$
$=\textsc{VCDL}(\bm{p},\bm{\theta})(\nabla u(1)-\nabla u(0))$
$\leq 2\textsc{VCDL}(\bm{p},\bm{\theta}).$
@@
$\textsc{CFDL}_{S}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S}(q_{i},\widehat{q}_{i})$ $=\lim_{j\to\infty}\frac{1}{T}\sum_{i=1}^{m}n_{i}\textsc{Breg}_{S_{j}}(q_{i},\widehat{q}_{i})$
$\leq 2\textsc{VCDL}(\bm{p},\bm{\theta}).$
Taking supremum over $S$ proves $\textsc{CDL}(\bm{p},\bm{\theta})\leq 2\textsc{VCDL}(\bm{p},\bm{\theta})$. ∎
@@ 4 CDL and Calibration Errors
    By Example 4.2, $\textsc{CDL}\leq 2\textsc{VCDL}\leq 4\epsilon$. We can find a Bregman divergence such that $\textsc{CFDL}\geq\frac{\epsilon}{2}$. Consider the scoring rule $S$ that has a V-Bregman divergence with kink $\mu=1-\frac{\epsilon}{2}$.
@@
    We can prove $\textsc{VCDL}=\Theta(1)$. Fix any V-Bregman divergence with kink $\mu$, consider the corresponding $\textsc{CFDL}_{\mu}$.
@@
    We know $\textsc{CDL}\leq 2\textsc{VCDL}\leq\frac{8}{T}$.
@@ 5.1 Attributing CDL to Bucket-wise Biases
$\textsc{VCDL}(\bm{p},\bm{\theta})\leq\frac{2m}{T}\mathcal{D}(\bm{p},\bm{\theta})+\frac{2\alpha}{\sqrt{T}}+2\beta+O(\frac{\alpha^{2}m\log m}{T}).$ (20)